What are the basic algebra equation rules for multiplication and division? I have the folllowing equation: CPI=EV/AC is the earned value equation, with the following information provided: .92= 172,500/AC
How do you isolate AC, when it is the denominator and get the other numbers on the other side for the answer?

Answer by Theo(13342)

.92= 172,500/AC

multiply both sides of the equation by AC to get:

.92*AC = 172,500

divide both sides of the equation by .92 to get:

AC = 172,500 / .92

solve for AC to get:

AC = 187,500

you move things around by adding or subtracting or multiplying or dividing both sides of the equation.

the equality remains the same.

if a = b, then a+7 = b+7

if a = b, then a-7 = b-7

if a = b, then a/7 = b/7

if a = b, then a*7 = b*7

if a = b, then log(a) = log(b)

if a = b, then a^7 = b^7

as long as a = b, you can always substitute a for b to get:

a = a

In your equation, you had .92 = 172500/AC

To move AC out of the denominator, all you had to do was multiply both sides of the equation by AC to get:

AC * .92 = (AC * 172500)/AC

The AC on the right side of the equation cancels out to get:

AC * .92 = 172500